We had an issue with the E-beam evaporator. The auto track that controls the valve function had an issue which got repaired but you will still notice a delay when opening the high vac. I believe the valve needs replaced. If it becomes a big problem notify staff.
We experienced problems with the CDA (clean compressed air) which started last Thursday. By midday the system dropped down to 80 psi. It affected the STS and Cannon stepper. Talked to facilities they had to rebuild the heads on the compressor. I believe this has been accomplished last Wednesday. Everything should be back to normal. If you experience problems let staff know.
We lost the dry pump on the ALD system. Jake is in the process of removing it and sending it out for repair. Looking for turnaround time to be 4 working days.
We have ordered the XeF2 and should have that shortly.
In the BiON on the Decktac 3, users were experiencing that the operating system would shut down and needed to be restarted. It would work for a short while and then shut down again. The back up got reinstalled into the system and so far it looks like the tool is functioning properly. Let staff know if you are experiencing any problems.
We are in the process of getting a new cylinder and will try and have it back up ASAP. We will keep you updated.
The STS & XeF2 are back up and running.